OpenStudy (anonymous):

What is the solution to the system of equations? 2x+y=-3 x-2y=-4

OpenStudy (anonymous):

@Katdetty can you help?

OpenStudy (anonymous):

Well first you solve one equation. Pick one of the equations and tell me which you would like to do first

OpenStudy (anonymous):

2x+y=-3

OpenStudy (anonymous):

Ok so, let's work on finding y. So since we are finding y. We put 2(0)+y = -3. And solve. So y=3

OpenStudy (anonymous):

Now plug y into the other equation

OpenStudy (anonymous):

x=-10? @Katdetty

OpenStudy (anonymous):

X-2(3)=-4. ==> X-6=-4 -4+6 = 2. So x=2

OpenStudy (anonymous):

And sorry my internet has been lagging a lot